# KoCoSa: Korean Context-aware Sarcasm Detection Dataset

### 22 Feb 2024 v1

@@ Abstract
We show that the sarcasm detection task relies deeply on the existence of sufficient context. We will release the dataset at https://anonymous.4open.science/r/KoCoSa-2372.
@@
| Dept. of Applied Statistics Dept. of Computer Science and Engineering |
| Dept. of Artificial Intelligence, Chung-Ang University |
@@ 1. Introduction
Meanwhile, a majority of datasets utilized for training sarcasm detection systems originate from English Twitter and Reddit comments (Moores and Mago, 2022). However, these datasets inherently diverge from daily dialogues that predominantly involve interactions among acquaintances, making it difficult to use the sarcasm detection system trained on these datasets for daily dialogues. In addition, sarcasm also reflects cultural differences that exhibit the nuances of each country’s culture. When annotated by annotators from different countries, these differences result in a degree of inconsistency (Liu et al., 2014; Joshi et al., 2016). Therefore, relying solely on translationese when investigating sarcasm detection methods in various languages is inappropriate, as the translationese may not capture the linguistic nuances of sarcasm. As a result, there is an imperative need for monolingual datasets, especially considering that research on English sarcasm detection has been notably more extensive compared to other languages (Rahma et al., 2023).
@@ 3.1. Collecting Source Dialogue
We use two primary Korean source dialogue corpora to construct KoCoSa: NIKL Messenger Corpus National Institute of Korean Language (2022a)(24.4%) and NIKL Online Text Message Corpus National Institute of Korean Language (2022b)(75.6%). The data formats of both corpora are nearly identical, with very minor differences. The subject matter of this source dialogue corpora pertains to daily dialogue. We find that in cases involving three or more participants, the conversation lacks consistency due to the unordered nature of utterance sequences in online messengers. Hence, we only use the dialogues that are made between two people for the dataset.
@@ 3.2.2. Generating New Dialogue from Situation
In Korean, ensuring a consistent level of politeness is of utmost importance. Hence, when a single speaker switches between formal and informal language, it results in a less natural dialogue flow. To resolve the issue, we aim to maintain the level of politeness of each speaker in original dialogue by obtaining the formality of the speaker from a pre-trained formal classifier111https://huggingface.co/j5ng/kcbert-formal-classifier. We incorporate this formal information into the input prompt to create a dialogue that preserves the formality of speakers in each utterance.
@@ 3.3.1. Filtering Toxic Dialogue
Sarcastic utterances occasionally take the form of offensive language, so it is necessary to filter out data that may cause potential harmfulness when the dataset is disclosed. We guarantee the dataset to be moderate through a total of two steps. First, We apply the moderation API of OpenAI222https://platform.openai.com/docs/guides/moderation to all data. Among 17,073 samples, we removed 23(0.0013%) dialogues that were labeled as "improper." Then, we additionally filtered five samples that included swear words through the manual data inspection process. For swear word filtering, we use a pre-defined frequently used toxic words list.
@@ 3.4.2. Annotation Guideline
Finally, we only use the dialogues where the labels of both "Context Abnormality Detection" and "Sarcasm Explanation Revision" are natural to construct the dataset. Among them, we use the data where the label of Sarcastic Response is also "Sarcasm" as the "Sarcasm" label in the dataset. For the "Non-sarcasm," we first include the data where the response is Non-Sarcasm/Abnormal and the context is Normal. Additionally, we use the dialogue in cases where the context is Normal but the response is Abnormal by removing the last sarcastic response which is a cause of abnormal. This approach of constructing non-sarcasm data helps address the issue of having data comprised solely of failed attempts at generating sarcasm dialogue.
@@ 4.2.1. Label Change
| Model | Balanced Acc. | Weighted-F1 | Precision-S | Recall-S | Precision-N | Recall-N |
| Zero/Few-shot | ||||||
@@ 5.1. Experimental Setup
We split KoCoSa into Train/Dev/Test in the proportions of about 8:1:1. To mitigate dataset bias, we exchange the name of speakers A and B for 50% of the dataset, as most of the sarcastic responses’ speaker in the dataset is A. The input data passed to the models consists of contexts, and sarcastic responses, where the output is the label for the sarcasm detection task. We do not use explanation data for developing sarcasm detection systems in our work, leaving it for future work.
@@
We use six metrics considering the imbalanced label distribution in the dataset; Balanced Accuracy, Weighted F1, Precision-[Sarcasm, Non-Sarcasm] and Recall-[Sarcasm, Non-Sarcasm].

### 22 Mar 2024 v2

@@ Abstract
We show that the sarcasm detection task relies deeply on the existence of sufficient context. We will release the dataset at https://github.com/Yu-billie/KoCoSa_sarcasm_detection.
@@
| Chung-Ang University, Seoul, Korea |
@@ 1. Introduction
Meanwhile, a majority of datasets utilized for training sarcasm detection systems originate from English Twitter and Reddit comments (Moores and Mago, 2022). However, these datasets inherently diverge from daily dialogues that predominantly involve interactions among acquaintances, making it difficult to use the sarcasm detection system trained on these datasets for daily dialogues. In addition, sarcasm also reflects cultural differences that exhibit the nuances of each country’s culture. When annotated by annotators from different countries, these differences result in a degree of inconsistency (Liu et al., 2014; Joshi et al., 2016). Therefore, relying solely on translationese111translations that are unnatural and overly literal. when investigating sarcasm detection methods in various languages is inappropriate, as the translationese may not capture the linguistic nuances of sarcasm. As a result, there is an imperative need for monolingual datasets, especially considering that research on English sarcasm detection has been notably more extensive compared to other languages (Rahma et al., 2023).
@@ 3.1. Collecting Source Dialogue
We use two primary Korean source dialogue corpora to construct KoCoSa: NIKL Messenger Corpus National Institute of Korean Language (2022a)(24.4%) and NIKL Online Text Message Corpus National Institute of Korean Language (2022b)(75.6%). The data formats of both corpora are nearly identical, with very minor differences. The subject matter of this source dialogue corpora pertains to daily dialogue. We find that in cases involving three or more participants, the dialogue lacks consistency due to the unordered nature of utterance sequences in online messengers. Hence, we only use the dialogues that are made between two people for the dataset.
@@ 3.2.2. Generating New Dialogue from Situation
In Korean, ensuring a consistent level of politeness is of utmost importance. Hence, when a single speaker switches between formal and informal language, it results in a less natural dialogue flow. To resolve the issue, we aim to maintain the level of politeness of each speaker in original dialogue by obtaining the formality of the speaker from a pre-trained formal classifier222https://huggingface.co/j5ng/kcbert-formal-classifier. We incorporate this formal information into the input prompt to create a dialogue that preserves the formality of speakers in each utterance.
@@ 3.3.1. Filtering Toxic Dialogue
Sarcastic utterances occasionally take the form of offensive language, so it is necessary to filter out data that may cause potential harmfulness when the dataset is disclosed. We guarantee the dataset to be moderate through a total of two steps. First, We apply the moderation API of OpenAI333https://platform.openai.com/docs/guides/moderation to all data. Among 17,073 samples, we removed 23(0.0013%) dialogues that were labeled as "improper." Then, we additionally filtered five samples that included swear words through the manual data inspection process. For swear word filtering, we use a pre-defined frequently used toxic words list.
@@ 3.4.2. Annotation Guideline
| Context(C) | Response(R) | Label | Final Format |
Normal |
Sarcasm |
Sarcasm | C;R |
Normal |
Non-Sarcasm |
Non-Sarcasm | C;R |
Normal |
Abnormal |
Non-Sarcasm | C |
Abnormal |
Deleted | ||
With the annotated data, we construct the dataset with the construction criteria described in Table 3. In cases where the context is Normal and the response is Abnormal, we use the dialogue after removing Response from the dialogue.
@@ 4.2.1. Label Change
| Model | BA | M-F1 | W-F1 | Precision-S | Recall-S | Precision-N | Recall-N |
| Zero/Few-shot | |||||||
@@ 5.1. Experimental Setup
We split KoCoSa into Train/Dev/Test in the proportions of about 8:1:1. Since most generated data start with speaker A, we exchange the names of speakers to mitigate the possible bias. By randomly selecting 50% of dataset, we switch the name of speaker A to speaker B, and the name of speaker B to speaker A.
The input data passed to the models consists of contexts, and sarcastic responses, where the output is the label for the sarcasm detection task. We do not use explanation data for developing sarcasm detection systems in our work, leaving it for future work.
@@
We use seven metrics considering the imbalanced label distribution in the dataset; Balanced Accuracy, Macro F1, Weighted F1, Precision-[Sarcasm, Non-Sarcasm] and Recall-[Sarcasm, Non-Sarcasm].
@@ 9. Language Resource References
| Change Type | Sarcasm –> Non-Sarcasm | |||||||
| Casual Dialogue |
context | ||||||
| A: 아니야, 직접 가서 사는 게 훨씬 편하잖아. (No, going there in person is much more convenient) | Response | |||||||
| Direct Criticism |
Context | ||||||
| A: 와우, 허세 정말 한가득이네. (Wow, that’s quite extravagant) | Response | |||||||
| Input Prompt |
You are Korean. You create natural Korean dialogues proficiently. Please consider the level of politeness. Level of politeness: A-{}, B-. |
Sarcasm: someone says something but means the opposite mockingly or ironically, often using tone and context to convey the real meaning. |
Level of politeness: A-반말(Informal), B-반말(Informal)) |
Intimacy: 4 |
A: 요리는 잘 돼가? (How’s the cooking going?) |
B: 응 지금까지는 순항 중이야. 하나만 빼고.(Yeah, it’s been smooth sailing so far. Except for one thing.) |
A: 뭐가 문제야? 잘 안되는게 있어? (What’s the problem? Is something not going well?) |
B: 계란 후라이가 조금 탔어. (The fried eggs are a little burnt.) |
Example 2. Situation: 자기계발-퇴근 후 자기계발을 위해 학원에 등록한 상황 (Self-improvement - Registering for classes after work for self-improvement) |
Level of politeness: A-존댓말(Formal), B-반말(Informal)) |
Intimacy: 3 |
A: 퇴근하고 뭐 하는거 있어요? (Do you do anything after work?) |
B: 아니 퇴근하면 힘들잖아. 그냥 집에 가서 쉬어야지. (No, I’m tired after work. I just go home and rest.) |
A: 저는 얼마 전에 영어학원 등록했어요. (I recently registered for an English class.) |
B: 아 진짜? 영어 공부 하려고?? 저번 달에는 중국어 공부할거라며? (Really? To study English?? Last month, you said you were going to study Chinese.) |
A: 중국어는 너무 어렵더라고요. 그래서 큰 돈주고 영어학원 다시 등록했어요. (Chinese turned out to be too difficult. So, I ended up registering for an English class, even if it cost a lot.) |
Sarcasm Explanation: 영어학원에 등록만 하고 가지 않을 것 같은 상대방의 행동을 긍정적인 기부를 하는 것처럼 비꼬아 말한다. (Sarcastically comments on the other person’s action of registering but not attending the English class as if it’s a generous donation.) |
Sarcastic response(B): 학원에 그렇게 기부를 많이 해도 되는거야? (Is it okay to make such generous donations to the academy?) |
Situation-{situation}, |
Level of politeness: A-{}, B-{} |
Intimacy: intimacy |
Generate Example: |
Appendix A Construction Details
A.1. Hyperparameters
We employ both GPT-4 and GPT-3.5 large language models (LLMs) for generation. For GPT-3.5, which is utilized to extract situations, we configure it with a maximum of 1000 tokens, a temperature of 1.0, top_p set to 1.0, and n set to 1. In the case of GPT-4, which is used for generating sarcastic dialogues, the settings include a maximum of 3000 tokens, a temperature of 1.1, top_p at 0.8, n set to 1, and a frequency penalty of 0.2. The full prompt including the two-shot examples is described in Table 12.
A.2. Annotation Details
A.2.1. Annotator Recruitment
We recruit annotators from the university’s online community. All were well-educated and native Koreans. We present the same 20 pieces of initially generated data (complete dialogue including both context and response) to 30 applicants. Annotation qualifications are granted to 10 applicants who show a match rate of 65% or more with the majority vote.
A.2.2. Annotator Compensation
All annotations were conducted in the web page constructed using streamlit444https://streamlit.io/. The total number of samples is 14,788 so each annotator labeled 1,420 to 1,904 samples. Figure 4 shows the annotation page used in the annotation process. We pay more than $10/hour, above the minimum wage, based on the hourly workload measured by the preliminary survey.
| Abnormal Context Case | Example |
| Unnatural Choice of Words |
A: 우리 아들, 요즘 왜 그렇게 책을 많이 읽어? (My son, why do you read so many books these days?)
| Inconsistent Opinion |
A: 아, 커피 너무 좋아. 아메리카노는 물맛 같고, 라테는 너무 달아서 맛없어. (Oh, I love coffee. Americano tastes like water and latte is too sweet, so I don’t like them)
| Absence of Dialogue |
Sarcasm explanation: 마지막 B의 말에 따르면, 실제로는 코로나 때문에 콘서트를 가지 못하는데 이것을 긍정적으로 비꼬아 말한다. (According to B’s last utterance, A is sarcastically making a positive mockery of the fact that B cannot attend concerts because of COVID-19.)
Appendix B Dataset Examples
As discussed in section 4.2, we aim to create natural dialogues where the final responses are sarcastic. However, LLMs generated unintentional data contrary to our expectations. Here, we show a few examples for the aforementioned data in 4.2.
B.1. Label Change Data
B.2. Abnormal Context Data
As described in Table 3, we fully utilized the generated data in various ways. Nevertheless, any instances where the context itself was labeled as abnormal were deleted. Here, we present examples of the removed data in Table 13.
Appendix C Experimental Details
C.1. Training KoCoSa
When training and on KoCoSa, we set the batch size to 16 and learning rate to 1e-5.
C.2. Instruction for GPT-3.5 and GPT-4
[…]